DIRC Operations Weekly Report

Period: Monday 2001/1/1 - Wednesday 2007/1/10

Oncall: Bradley Wogsland
Fast Monitoring: Bradley Wogsland, Alejandro Perez

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20070110.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20070110 (empty)

Contents


BaBar-PEP news

PEP peak lumi: 0 x 1033 (previous week: NA)

BaBar integrated lumi: 0 fb-1 (previous week: NA)

Run 6 PEP delivery: 0 fb-1

Run 6 total delivery: 0 fb-1 Run 6 total BaBar recorded: 0 fb-1 including 0 fb-1 off-peak data

PEP activities this week:

  • Run 6 startup tuning. Beams soon.

    Foreseen work opportunities (PEP down periods):

  • Day shift Jan 17 & 31.

    BaBar activities/issues this week:

  • Taking cosmics since the weekend

    DIRC activities

    See DIRC logbook for details

    Backgrounds

    NA

    Scheduled activities

    Pending Work

    Manpower situation

    Bradley, Jose and Alejandro rotating backup & on-call commissioning duties, with Alejandro working doubletime.

    Links